OpenStudy (anonymous):

What is the magnitude of the force on an electron moving with a velocity of 6.33 × 106 m/s at right angles to a 2.53 × 10-2 T magnetic field? 2.56 × 10-14 N 2.56 × 10-11 N 2.56 × 10-8 N 2.56 × 10-5 N 2.56 × 10-2 N

OpenStudy (ujjwal):

F= e(V x B) = eVB sin theta...

OpenStudy (anonymous):

e = ?? theta =?? B=2.53 × 10-2 T ??

OpenStudy (ujjwal):

e is electronic charge. theta is angle between velocity of electron and magnetic field.. B is magnetic field.

OpenStudy (anonymous):

theta will be 180 ? e =1.602×10^-19 ????

OpenStudy (esraa):

2.56 x 10^-14

OpenStudy (anonymous):

1 ans is correct